# Understanding User Mental Models in AI-Driven Code Completion Tools: Insights from an Elicitation Study

### v1

@@ Abstract
Integrated Development Environments increasingly implement AI-powered code completion tools (CCTs), which promise to enhance developer efficiency, accuracy, and productivity. However, interaction challenges with CCTs persist, mainly due to mismatches between developers’ mental models and the unpredictable behavior of AI-generated suggestions. This is an aspect underexplored in the literature. To address this gap, we conducted an elicitation study with 56 developers using focus groups, to elicit their mental models when interacting with CCTs. The study findings provide actionable insights for designing human-centered CCTs that align with user expectations, enhance satisfaction and productivity, and foster trust in AI-powered development tools. To demonstrate the feasibility of these guidelines, we also developed ATHENA, a proof-of-concept CCT that dynamically adapts to developers’ coding preferences and environments, ensuring seamless integration into diverse workflows.
@@
journal: Anonimyzed journal
[…]
orcid]https://orcid.org/0000-0001-9894-2116 orcid]https://orcid.org/0000-0002-9536-3087 orcid]https://orcid.org/0000-0003-2730-7697 orcid]https://orcid.org/0000-0001-5181-7115 orcid]https://orcid.org/0000-0002-1421-3686 orcid]https://orcid.org/0000-0003-1561-70736
@@ 1 Introduction
Integrated Development Environments support developers with valuable tools to write and debug code in current software development practices. In many IDEs, code suggestion by automated systems has become a key element. Employing ever-developing Artificial Intelligence (AI) capabilities, these Code Completion Tools intend to boost developer efficiency, accuracy, and productivity by cutting down on the manual input needed for repetitive code segments and speeding up the entire coding task. CCTs powered by AI, like GitHub Copilot and JetBrains AI, indicate a movement towards advanced and informed programming assistance.
@@
When dealing with CCTs, while potentially benefitting from increased efficiency, users are exposed to a wide range of risks strictly related to automation. For example, one of the main risks in the domain of code completion is deskilling [1], i.e., the risk of losing personal abilities in the automated task (in this case, code writing). Similarly, and as already highlighted by previous studies [2], CCTs may increase the risk of generation of unsecured code. In response to the challenges that generally affect interaction with AI systems, as in the case of CCTs, a new field of study has emerged in recent years, lying at the intersection of Human-Computer Interaction (HCI) and AI, commonly known as Human-Centered Artificial Intelligence (HCAI) [3, 4]. HCAI proposes to develop systems that are designed, developed, and evaluated by involving users in the process, aiming to increase performances and satisfaction in specified tasks [5].
While the technical aspects of AI’s role in code generation are well-researched (for example, algorithms and AI models used by such tools [6, 7]), HCAI systems for code completions are still underresearched, exposing developers (and their software) to various risks. In fact, there is little indication on the interaction of developers with these tools, for example, how a system function is represented in a user’s mental model, which crucially affects the efficiency and success of using these tools. Unlike conventional code suggestion systems, AI tools frequently show unpredictable and changing behavior that might not match users’ mental models. When interpreting the advice these tools give or when choosing to accept and employ them, developers may experience challenges. Although some studies explored coder’s expectations when dealing with CCTs [2, 8], the understanding of the cognitive and interaction dynamics that developers face upon adopting CCTs is limited. Thus, the necessity for thoroughly investigating how users engage with CCTs has become crucial. Understanding developers’ mental models in using such tools, useful in a human-centered approach, can lead to developing IDEs that align with user requirements, foster user satisfaction, and simultaneously improve coding productivity while promoting their adoption [2].
This study addresses the lack of understanding of the user’s mental model in the usage of CCTs by reporting an elicitation study carried out as 8 focus groups with a total of 56 participants. Driven by the limitations identified during the analysis of state of the art, the study aims to understand specific aspects of the interaction with CCTs: where to show the suggestions, when the suggestion should be triggered during the interaction, how to show the suggested code, which type of code should be suggested, and personalization of the suggestion. Moreover, we also focus on various aspects of the interaction with the explanation of the tool’s suggestion: where to show the explanation, when the explanation should be shown, and what should be explained. Insights gained from this analysis will help dictate strategies for creating user-friendly interaction approaches in IDEs while increasing cohesiveness among developers and AI technologies in software development.
@@
The manuscript is structured as follows. Section 2 presents the rationale and background of our study, providing a survey of existing CCTs and their interaction modalities. Section 3 presents the methodology of our study, detailing the focus groups. Section 4 reports the findings of our elicitation study. Section 5 discusses the findings, presenting developers’ preferences and their mental model of human-centered AI-based CCTs. Section 6 illustrates a first working prototype that implements the elicited mental model. Finally, Section 7 concludes the article, presenting its limitations and potential future work.
@@ 2 Rationale and Background
With the rise of AI in software development environments, AI-driven CCTs are quickly revolutionizing developers’ approach toward programming tasks. Real-time assistance is provided through these tools with suggestions for code snippets, auto-completion of code syntax, and predicting coding patterns while improving productivity, lowering the cognitive load, and streamlining the coding process. Although the implementation and usability of CCTs very much rely on the developer’s expectations, understanding, and workflows, the success of these tools largely relies on how closely they align with what developers are already using, what they expect to use, their skills, and their needs. Achieving this alignment mainly hinges on users’ (internal) mental models or internal representations of how the system works.
A mental model is generally a mental picture users have of how a certain system works [9, 10]. It includes a user’s assumptions, beliefs, and expectations while interacting with a system that will lead to their actions, interactions, and problem-solving behavior. Mental models also change over time, might contain errors, and are constrained by the user’s prior experiences or technical background. Multiple mental models of a single object may be held by one person to reflect its many functions, and two persons will not always have the same mental model. If mental models are accurate and considered in the design of a system, users can anticipate system responses, feel confident during the interaction, and adapt quickly to new features. On the other hand, gaps between the users’ mental model of the system’s functionality and its actual behavior can lead users to confusion, frustration, and errors, which will reduce their productivity and satisfaction [11].
Mental models are crucial in the context of CCTs supported by AI. These tools employ Machine Learning (ML) algorithms or Large Language Models, thus relying on probabilistic computations rather than being deterministic. For example, AI-driven suggestions may use aggregated pattern data, which are not directly readable by users. Thus, the suggestions produced might be contextually nonrelevant or impossible to interpret. If a developer’s mental model does not match the underlying mechanisms of an AI tool, users might struggle to trust the tool, interpret its suggestions accurately, or recognize when and how to use it effectively.
The need to understand mental models in the domain of AI-based CCTs is multifaceted, and different aspects can benefit from a deep understanding of users’ mental models. The first is interaction with CCTs since users might predict the tool’s behavior and know when a suggestion will be useful and when it could mislead. For instance, a tool designed considering users’ mental models might allow developers to know when the AI might make contextually relevant suggestions and when it can blunder, and so on, to keep the coding flow in hand and avoid wasting time for distraction. Another aspect that can benefit regards trust. CCTs can be largely obscure to those operating outside the normal programming logic. It has been proven that transparency and interpretability are essential for establishing user trust, and understanding the user’s mental model is essential to making users comfortable when adopting the tools [12]. This helps create trust in a tool by aligning design features with users’ mental models, making developers more likely to adopt it into their daily workflow. Another aspect that can benefit from the mental model is cognitive load: CCTs are meant to reduce repetitive tasks and cognitive load; thus, users can concentrate on more difficult problem-solving jobs. However, if users cannot intuitively understand how or why these suggestions are being made, the cognitive load can actually rise rather than fall. Refining these tools to become less disjointed in relation to user mental models means we can improve the flow of work, minimize interruptions, and ultimately help users be more productive. The last aspect is the possible support to learning: especially for novice programmers, developers can recognize common coding patterns and problem-solving techniques, enhancing both tool competency and coding proficiency. The overall improvements in systems that stem from following users’ mental models throughout the design are also a potent driver for increasing systems’ adoption and acceptability [2]. A recent study by Wang et al. highlights that 54% of the developers recruited for their study are eager for better CCTs [8].
Exploring the users’ mental models is a complex process since these models are typically abstract, partially formed, or even subconscious. Without elicitation studies, bridging this gap is an arduous task, and they are the ones that prescribe structured methodologies to help yield, analyze, and interpret users’ cognitive models regarding system functionality. The techniques for eliciting mental models include design workshops [13], interviews [14], concept mapping [13], think-aloud protocols [15], and scenario-based design [16]. These techniques allow researchers to design intricate representations of users’ mental models of system behavior, what they expect the system to do a priori, and how they can expect to interact with, for example, CCTs. Such insights inform tool design and can help bolster our understanding of HCI within AI-augmented coding domains towards supporting frameworks and guidelines for improved user experiences in such environments and beyond.
In the next sub-sections, the state of the art concerning two important aspects of this study will be discussed, namely (i) the interaction currently offered by existing CCTs, and (ii) studies in the field of HCI on CCTs.
2.1 AI Code Completion Tools
The analysis of the remaining 13 CCTs providing a User Interface (UI) was guided by two main aspects: the code suggestion itself and the explanation, if any, of why that code was suggested. After familiarizing ourselves with the CCTs, we analyzed users’ interaction with them across different dimensions. To simplify the analysis and the reporting of the results, we used the 5W model, which is adopted in various fields, such as journalism and customer analysis, and more generally in problem-solving, to analyze the complete story of a fact. Each dimension of the interaction with the tool was addressed in terms of the question of the 5W model. In particular, the analysis of the tools with respect to the code suggestion was done by answering these questions:
  1. 1)
    Where to show the code suggestion in the UI?
  2. 2)
    When should the code suggestions be activated during the interaction?
  3. 3)
    How to show the suggested code?
  4. 4)
    What should be suggested?
  5. 5)
    How should the suggestion be customized?
  1. 6)
    Where to show the explanation in the UI
  2. 7)
    When to show the explanation during the interaction
  3. 8)
    What to explain.
A summary of the complete set of such models is depicted in Table 1.
Code completion Explanation
What When How to show How to customize Where What When Where
Intellicode [17] lines proactive grey text curent file inline - - -
Codex [18] lines manual text chat history right side answer prompt chatbot
Copilot [19] lines proactive grey text current file inline answer prompt chatbot
TravTrans [20] token proactive dropdown current file inline - - -
StarCoder [21] lines proactive grey text current file inline - - -
Codegeex [22] lines proactive grey text current file inline answer, description prompt, click chatbot
CodeWhisperer [23] lines hybrid grey text current file inline source click sidebar
Codeium [24] lines proactive grey text current file inline improvements, reasons click chatbot
Replit [25] lines proactive grey text current file inline description click pointer
Cody [26] lines proactive grey text repository inline answer click, prompt chatbot
Tabnine [27] lines proactive grey text current file inline answer prompt chatbot
Xcode [28] lines proactive grey text current file inline - - -
Blackbox [29] lines proactive grey text current file inline improvement click sidebar
2.1.1 Code suggestions
An important characteristic affecting the interaction with CCTs is the timing to provide suggestions (when). The analysis of the existing tools revealed two main approaches: proactive and manual. Proactive CCTs provide suggestions at a certain time without requiring a user action to trigger the completion. This type of design can be either implemented in an intrusive manner (by suggesting a completion at any time) or discreetly (by generating completions only when the context is sufficient). The manual approach, in contrast, requires the user to perform a specific action to visualize the completion. However, manual activation of CCTs is not very popular, as it is adopted in only two tools. A hybrid timing approach is also implemented, even if only in the case of CodeWhisperer, as completions are proactive but can be forced through a shortcut [23].
Another aspect is the content of CCTs suggestions (what is provided as a suggestion?). Two main modalities emerged from our analysis: single token suggestions (less popular, as it was proposed only with the TravTrans prototype [20]) and multiple line completions, which are widely adopted. While the first solution is quite simple, the second one results in a feature that is hard to control and assess since the length of the suggestions may vary depending on the context and LLM characteristics. As a matter of fact, multi-line suggestion length can range from short code blocks (e.g., for loops, switches, if-else) to entire functions or classes.
The suggestion’s location (where) in the UI is also worth exploring. A predominant approach is to put completions directly inside the code (inline suggestions). The other solution instead consists of providing the suggestion in a separate space with respect to the coding area. Codex [18] is the only tool that adopts the latter modality.
It is useful to discuss how to show code completions that appear before they are accepted and integrated into the live code. A consolidated approach is to display the completed code in gray, allowing users to easily distinguish it from the live code, which is typically shown in white (or other colors, depending on the users’ preferred color scheme), before accepting the proposals of the CCT. However, another method has emerged: displaying multiple possible completions in a dropdown menu. This latter approach may be particularly convenient for single-token suggestions or very short completions, as demonstrated by TravTrans.
Finally, it is worth investigating the customizability of CCTs, as the relevance and helpfulness of suggestions strongly depend on the type of user receiving them (how to customize). Context-aware suggestions may consider the user expertise, the specific formatting/adherence to coding standards, and the functional and non-functional requirements of specific software development scenarios to provide the most accurate code snippets possible. The strategies adopted for this aim result in three categories: personalization based on the context of the current file, personalization based on a code repository, and personalization based on chat history. The first category is the most popular, while Cody [26] is the only tool that managed to include a code repository for more relevant completions. Nevertheless, this last approach seems to fail for big repositories. Codex provides more aware suggestions by processing the conversation between the programmer and the chatbot as context.
2.1.2 Explanations of the suggested code
Some of the CCTs analyzed do not offer any explanatory feature. By contrast, the remaining tools use various methods to provide developers with explanations as they write new code.
The location in the UI (where) of code explanations is an important aspect of CCTs, as it influences other interaction characteristics. Three main approaches are observed: explanations are either delivered inside a chatbot, or shown in a sidebar, or finally displayed inside the coding space. Tabnine [27], Cody [26], CodeGeeX [22], Copilot [30], Codeium [24], and Codex [18] implement the first approach, providing chat-based explanations. In contrast, CodeWhisperer [23] and Blackbox [29] report code explanations in a sidebar. In this case, the user is not allowed of submitting prompts for additional details, but such an approach delivers code explanations quicker, as the user does not have to think about a prompt in order to receive explanations. Finally, Replit [25] follows a unique approach, showing explanations inside a tooltip at the mouse pointer’s location (which means the explanations appear within the coding space).
@@ 2.2 Usability studies on code completion tools
Another aspect investigated by HCI researchers focused on the users’ behavior while adopting CCTs. From this study, two main modalities emerged [32]: in “acceleration” mode, the programmer is aware of their next steps and utilizes the tool to speed up the coding process. In contrast, in the “exploration” mode, the programmer is uncertain about how to proceed and employs the tool to investigate his options. Acceleration mode is usually preferred, as developers often use the tools to complete repetitive code that cannot be copy-pasted [33], and for recalling syntax they don’t remember [34]. As a result, the more developers accept CCTs suggestions, the more they perceive themselves as productive [35].
Liang et al. conducted a study to understand the usability of CCTs (such as GitHub Copilot and Tabnine) by utilizing a structured survey distributed to 410 GitHub users who had engaged with CCTs [33]. Questions covered participants’ motivations for using or not using AI assistants, the usability issues they encountered, and the strategies they used to optimize the tool’s output. The survey revealed that developers are most motivated to use AI programming assistants because they help developers reduce keystrokes, finish programming tasks quickly, and recall syntax. They also found that developers often do not use CCTs because these tools do not generate code that addresses certain functional or non-functional requirements and because developers have trouble controlling the tool to generate the desired output.
Another important requirement that emerged from user studies is the need for CCTs to learn from feedback and adapt their behavior to developers’ styles and project needs. For instance, in [36], it emerged that programmers would like to customize the shortcuts and to set suggestions length and frequency [33]. The literature also highlights the need for explanations (such as reporting the source or linking to documentation) for additional context on the generated code [33, 34].
The proactivity of CCTs has also been investigated, with research encouraging its implementation [37]. Vaithilingam et al. suggested that automatic inline gray-text suggestions, as opposed to traditional lightbulb interfaces, significantly improve discoverability and usability of CCTs [38]. The authors established five design principles to guide future AI-driven code suggestion interfaces:

### v2

@@ Abstract
Integrated Development Environments increasingly implement AI-powered code completion tools (CCTs), which promise to enhance developer efficiency, accuracy, and productivity. However, interaction challenges with CCTs persist, mainly due to mismatches between developers’ mental models and the unpredictable behavior of AI-generated suggestions, which is an aspect underexplored in the literature. We conducted an elicitation study with 56 developers using co-design workshops to elicit their mental models when interacting with CCTs. Different important findings that might drive the interaction design with CCTs emerged. For example, developers expressed diverse preferences on when and how code suggestions should be triggered (proactive, manual, hybrid), where and how they are displayed (inline, sidebar, popup, chatbot), as well as the level of detail. It also emerged that developers need to be supported by customization of activation timing, display modality, suggestion granularity, and explanation content, to better fit the CCT to their preferences. To demonstrate the feasibility of these and the other guidelines that emerged during the study, we developed ATHENA, a proof-of-concept CCT that dynamically adapts to developers’ coding preferences and environments, ensuring seamless integration into diverse workflows.
@@
journal: Anonymized journal
orcid]https://orcid.org/$0$-0001-9894-2116 orcid]https://orcid.org/$0$-0002-9536-3087 orcid]https://orcid.org/$0$-0003-2730-7697 orcid]https://orcid.org/$0$-0001-5181-7115 orcid]https://orcid.org/$0$-0002-1421-3686 orcid]https://orcid.org/$0$-0003-1561-70736
@@ 1 Introduction
Integrated Development Environments support developers with valuable tools to write and debug code in current software development practices. In many IDEs, code suggestion by automated systems has become a key element. Employing ever-developing Artificial Intelligence (AI) capabilities, these Code Completion Tools intend to boost developer efficiency, accuracy, and productivity by cutting down on the manual input needed for repetitive code segments and speeding up the entire coding task (Weber et al., 2024; Cui et al., 2024; Peng et al., 2023). CCTs powered by AI, like GitHub Copilot and JetBrains AI, indicate a movement towards advanced and informed programming assistance.
@@
When dealing with CCTs, while potentially benefitting from increased efficiency, users are exposed to a wide range of risks strictly related to automation. For example, one of the main risks in the domain of code completion is deskilling (Sambasivan and Veeraraghavan, 2022), i.e., the risk of losing personal abilities in the automated task (in this case, code writing). Similarly, and as already highlighted by previous studies (Sergeyuk et al., 2025), CCTs may increase the risk of generating unsecured code. In response to the challenges that generally affect interaction with AI systems, as in the case of CCTs, a new field of study has emerged in recent years, lying at the intersection of Human-Computer Interaction (HCI) and AI, commonly known as Human-Centered Artificial Intelligence (HCAI) (Shneiderman, 2022; Xu, 2019). HCAI proposes to develop systems that are designed, developed, and evaluated by involving users in the process, aiming to increase performances and satisfaction in specified tasks (Desolda et al., 2024).
[…]
This study addresses the lack of understanding of the user’s mental model in the usage of CCTs by reporting an elicitation study carried out as 8 co-design workshops with a total of 56 participants. Driven by the limitations identified during the analysis of the state of the art, the study aims to answer the following research questions:
  1. RQ1:
    How do developers mentally model their interaction with AI-driven CCTs?
  2. RQ2:
    What are developers’ expectations regarding the delivery of explanations alongside AI-generated code suggestions?
  3. RQ3:
    How do developers perceive and articulate their needs for personalization in CCTs?
  4. RQ4:
    What human-centered design principles emerge for aligning CCTs with developer mental models?
Insights gained from this analysis will help dictate strategies for creating user-friendly interaction approaches in IDEs while increasing cohesiveness among developers and AI technologies in software development.
1.1 Contributions
This work advances the understanding and design of AI-driven Code Completion Tools (CCTs) from a Human-Centered AI perspective by making the following key contributions:
  • Exploratory Survey of CCTs: An exploratory survey has been conducted to identify and analyze AI-assisted CCTs, their interaction modalities, and explanation features. We contributed in this way by establishing the landscape of available approaches. Moreover, this activity informed the design of our elicitation study, ensuring that workshops prompts addressed both prevalent and underexplored aspects of developers interacting with CCTs.
  • Empirical mapping of developer mental models: we contribute to the state of the art by proposing the first qualitative elicitation of mental models of CCTs, to explore how developers conceptualize, interact with, and personalize CCTs. Our analysis revealed the multifaceted preferences regarding when code suggestions should appear (manual, proactive, hybrid), where and how they should be displayed (inline, sidebar, popup, chatbot), the granularity and content of suggestions, and the need for contextual, purpose-driven explanations.
  • Actionable design guidelines for human-centered CCTs: study findings have been distilled into concrete design guidelines, which emphasize that no single interaction model can fit all users or tasks. Indeed, users need highly customizable and adaptive code and explanation suggestions, including per-user control over activation mode, visualization, suggestion scope, explanation detail, and integration with individual and project coding styles.
  • Prototype demonstration of personalizable CCTs: a prototype of CCT called ATHENA has been developed to demonstrate the feasibility of the proposed guidelines, as well as to provide a platform for future comparative studies on developer-AI co-adaptation.
Together, these contributions provide a significant contribution to the HCI literature on AI development tools, providing evidence and solutions for designing trustworthy, effective, and personally relevant CCTs.
1.2 Organization
@@
The manuscript is structured as follows. Section˜2 presents the rationale and background of our study, providing a survey of existing CCTs and their interaction modalities. Section˜4 presents the methodology of our study, detailing the co-design workshops. Section˜5 reports the findings of our elicitation study. Section˜6 discusses the findings, presenting developers’ preferences and their mental model of human-centered AI-based CCTs. Section˜7 illustrates a first working prototype that implements the elicited mental model. Section˜8 discusses the elements of the study that could threaten its validity and describes the measures taken to mitigate them. Finally, Section˜9 concludes the article, presenting its limitations and potential future work.
@@ 2 Rationale and Background
AI-driven CCTs are influencing developers’ approach toward programming tasks (Alenezi and Akour, 2025; Martinović and Rozić, 2024). Real-time assistance is provided through these tools with suggestions for code snippets, auto-completion of code syntax, and predicting coding patterns while improving productivity, lowering the cognitive load, and streamlining the coding process. Although the implementation and usability of CCTs very much rely on the developer’s expectations, understanding, and workflows, the success of these tools largely relies on how closely they align with what developers are already using, what they expect to use, their skills, and their needs. Achieving this alignment mainly hinges on users’ mental models, i.e., internal representations of how the system works.
2.1 Human-centered AI and code completion tools
The rapid growth of AI in recent years has significantly increased its adoption and impact in research and business activities. However, AI’s pervasiveness has exacerbated concerns over existing flaws. In particular, biases and lack of explainability endanger the users of the AI models, which often do not consider the human element (Shneiderman, 2020). A report by the National Transportation Safety Board (2017) regarding a deadly crash of an autonomous Tesla car stated the following: <<automation “because we can” does not necessarily make the human-automation system work better. […] This crash is an example of what can happen when automation is introduced “because we can” without adequate consideration of the human element>>. To mitigate these issues, the novel field of HCAI is gaining traction (Shneiderman, 2022). One of the main goals of HCAI is to produce AI systems that are Reliable, Safe, and Trustworthy (Shneiderman, 2020). For this purpose, AI systems should provide a high level of computer automation while guaranteeing a high level of human control when desired—i.e., AI systems should not automate tasks but amplify, augment, empower, and enhance their users’ skills.
In the context of CCTs, systems can be designed either by aiming at full automation or at full augmentation (or anywhere in between). On one hand, a full-automation approach can provide systems that are extremely efficient, they are associated with several risks. In addition to an increase in commission and omission errors (Cummings, 2004), full-automation CCTs increase the risk of deskilling (Sambasivan and Veeraraghavan, 2022). On the other hand, previous user studies highlighted that using a full-augmentation approach “by default” may lead to systems that are not on par with users’ expectations, thus making them inefficient and ineffective (Esposito et al., 2024). In this context, our study emphasizes the need to design for augmentation, while still aligning CCTs with users’ mental models, thus ensuring that developers retain full control over their tools and allowing them to balance the two approaches.
@@ 2.2 Usability studies on code completion tools
Another aspect investigated by HCI researchers focused on the users’ behavior while adopting CCTs. From this study, two main modalities emerged (Prather et al., 2024): in “acceleration” mode, the programmer is aware of their next steps and utilizes the tool to speed up the coding process. In contrast, in the “exploration” mode, the programmer is uncertain about how to proceed and employs the tool to investigate his options. Acceleration mode is usually preferred, as developers often use the tools to complete repetitive code that cannot be copy-pasted (Liang et al., 2024), and for recalling syntax they don’t remember (Xu et al., 2022). As a result, the more developers accept CCTs suggestions, the more they perceive themselves as productive (Bird et al., 2022).
Liang et al. (2024) conducted a study to understand the usability of CCTs (such as GitHub Copilot and Tabnine) by utilizing a structured survey distributed to 410 GitHub users who had engaged with CCTs. Questions covered participants’ motivations for using or not using AI assistants, the usability issues they encountered, and the strategies they used to optimize the tool’s output. The survey revealed that developers are most motivated to use AI programming assistants because they help developers reduce keystrokes, finish programming tasks quickly, and recall syntax. They also found that developers often do not use CCTs because these tools do not generate code that addresses certain functional or non-functional requirements, and because developers have trouble controlling the tool to generate the desired output.
Another important requirement that emerged from user studies is the need for CCTs to learn from feedback and adapt their behavior to developers’ styles and project needs. For instance, Zhang et al. (2023) highlight that programmers would like to customize the shortcuts and to set suggestions length and frequency. The literature also highlights the need for explanations (such as reporting the source or linking to documentation) for additional context on the generated code (Liang et al., 2024; Xu et al., 2022).
The proactivity of CCTs has also been investigated, with research encouraging its implementation (Sergeyuk et al., 2024). Vaithilingam et al. (2023) suggested that automatic inline gray-text suggestions, as opposed to traditional lightbulb interfaces, significantly improve discoverability and usability of CCTs. The authors established five design principles to guide future AI-driven code suggestion interfaces: